# Performance-lossless Black-box Model Watermarking

### v1

@@ 6. Conclusion
Existing model watermarking work in black-box scenarios mainly relies on the idea of training-based backdoor, resulting in distortion to the original task performance. In this paper, we first propose a branch backdoor-based model watermarking protocol to protect model intellectual property, and give a performance-lossless construction based the message authentication scheme. The performance-lossless property of the proposed is reduced to the EUF-CMA security of MAC. Then, taking the language generation task as an example and combined with existing generative text watermarking, we instantiate a performance-lossless model watermarking which is resistant against potential interference attacks.
7. Acknowledgement

### v2

@@ VI Conclusion
Existing model watermarking work in black-box scenarios relies mainly on the idea of training-based backdoors, resulting in distortion of the primary task performance. In this paper, we first propose a branch backdoor-based model watermarking protocol to protect model intellectual property and propose a performance-lossless construction based the message authentication scheme. The performance-lossless property of the proposed BranchWM is reduced to the EUF-CMA security of MAC. Then, we analyze the possible attacks during the forensic process and instantiate a secure and performance-lossless watermarking for language generation model.
Acknowledgments
@@
Biography
[Uncaptioned image] Na Zhao received his B.S. degree in 2017 from Zhengzhou University (ZZU). Currently, she is working toward the Ph.D. degree in the University of Science and Technology of China. Her research interests include information hiding, and deep learning security.
[Uncaptioned image] Kejiang Chen received his B.S. degree in 2015 from Shanghai University (SHU) and a Ph.D. degree in 2020 from the University of Science and Technology of China (USTC). Currently, he is an associate researcher at the University of Science and Technology of China. His research interests include information hiding, and deep learning security.
[Uncaptioned image] Weiming Zhang received his M.S. degree and Ph.D. degree in 2002 and 2005, respectively, from the Zhengzhou Information Science and Technology Institute, P.R. China. Currently, he is a professor with the School of Information Science and Technology, University of Science and Technology of China. His research interests include information hiding and multimedia security.
[Uncaptioned image] Nenghai Yu received his B.S. degree in 1987 from Nanjing University of Posts and Telecommunications, an M.E. degree in 1992 from Tsinghua University and a Ph.D. degree in 2004 from the University of Science and Technology of China, where he is currently a professor. His research interests include multimedia security, multimedia information retrieval, video processing and information hiding.